# Efficient Pyramid Channel Attention Network for Pathological Myopia Recognition with Pretraining-and-Finetuning

### v2

@@ 4.1 Datasets
PM-Fundus: To the best of our knowledge, the most existing datasets for PM recognition are private, lacking public PM recognition datasets. In order to prompt the development of PM recognition task, we collect fundus images of PM from publicly available PLAM dataset, ODIR dataset and RFMiD dataset pachade2021retinal to construct a new dataset named PM-fundus, which will be released as a benchmark for automatic PM recognition task. The PM-fundus dataset contains 818 images of PM and 1636 images of normal. We split the PM-Fundus dataset into two disjoint sets: training and setting. The training set contains 572 images of PM and 1145 images of normal, and the testing set contains 246 images of PM and 491 images of normal, which will be released in the future.
@@
LAG: It is a publicly available fundus image dataset of glaucoma li2019large , which was collected by Beijing Tongren Hospital. The dataset has 4,854 images: 1,711 images of glaucoma and 3,143 images of non-glaucoma. We also split it into three disjoint subsets: training (2,911), validation (971), and testing (972).
@@ 4.2 Baselines
In this paper, we use the following attention methods to verify the effectiveness of our EPCA module: SE, SRM, ECA, SPA, external attention (EA) guo2022beyond , coordination attention (CA) hou2021coordinate , and convolutional bottleneck attention module (CBAM) woo2018cbam . In addition, recent deep neural networks, such as vision transformer (ViT) dosovitskiy2020image , Res-MLP  touvron2021resmlp , swin transformer (Swin-T)  liu2021swin and pyramid vision transformer (PVT)  wang2021pvtv2 , and deep neural networks specifically designed for medical image analysis like DR-Net guo2020dense , MSCA-Net fu2023rmca , MSA-Net 9363019 , FIT-Net chen2023fit , RIR-Net ZHANG2022102499 , which are also adopted for comparison.
@@ 4.3 Implementation details
In this paper, we implement our EPCA-Net and other comparable deep neural networks with the Pytorch tool by using pre-trained backbone networks (such as ResNet18 and ResNet50) with traditional finetuning paradigm. We run all experiments on a server with an NVIDIA TITAN V GPU (11GB RAM). The SGD optimizer is utilized to optimize the deep networks with default settings (a weight decay of 1e-5 and a momentum of 0.9). We set the batch size and training epochs to 32 and 100, respectively. The initial learning rate (lr) is set to 0.0015, and the cosine annealing warm restart method loshchilov2016sgdr is applied to adjust lr. The classical data augmentation methods like flipping and rotation ($-10^{\circ}$-$10^{\circ}$) are used for data augmentation in the training.
@@ 4.4 Evaluation Metrics
Following existing work sun2023deep , we adopt the five evaluation metrics to verify the general performance of methods: accuracy (ACC), sensitivity (Sen), F1, kappa coefficient value, and the area under the ROC curve (AUC). In particular, AUC is applied to evaluate the binary classification performance of methods.
@@ 5.1.1 The impact of receptive field size number selection in pyramid pooling
As shown in Eq. 1, the receptive field size number selection in the pyramid pooling operator has a significant impact on the EPCA module. In this part, we train them with EPCA by using Linear-SCFM implementation on the Fundus-PM dataset and the Isee-PA dataset. Here, we adopt the pre-trained ResNet18 and ResNet50 as the baselines. Table 1 lists the classification results of our EPCA with six different receptive field size selection settings. We see that multi-scale context features obtained by spatial average pooling operations with different receptive field sizes clearly affect the performance of EPCA, demonstrating that multi-scale context features have varying significance levels. When pyramid pooling with multiple receptive filed size setting 1,3, our EPCA achieves the best results on two datasets, which is used for the following experiments. Moreover, EPCA based on ResNet50 performs better than it based on ResNet18; thus, ResNet50 is adopted as the baseline for the following ablation experiments.
Table 1: Performance comparisons of receptive field size number selection in pyramid pooling based on the EPCA module
Receptive field size setting PM-Fundus Isee-PA
ResNet18 ResNet50 ResNet18 ResNet50
$\left\{1\right\}$ 96.07 96.47 74.27 74.41
$\left\{1,2\right\}$ 96.61 96.20 74.83 76.63
$\left\{1,3\right\}$ 96.88 97.56 76.50 77.19
$\left\{1,4\right\}$ 96.07 96.93 76.63 75.80
$\left\{1,2,3\right\}$ 95.93 96.47 74.69 75.94
$\left\{1,2,4\right\}$ 95.79 95.93 76.36 76.36
@@ 5.1.2 The impact of different SCFM implementations
Fig. 5 presents the classification results of the EPCA module with four different SCFM implementations on the Fundus-PM (green color) dataset and the Isee-PA (blue color) dataset. It can be seen that EPCA module with Linear-SCFM and Hierarchical-SCFM achieves similar performance and outperforms it with Dropout-SCFM and Parallel-SCFM. The following reasons can explain the results: 1) The Dropout technique randomly selects context features, which makes it challenging to evaluate whether the selected context features are helpful while the rest of the context features are useless. 2) Although multi-scale context features with varying significance levels are extracted from different receptive sizes based on spatial average pooling, every receptive filed size produces attention weights in the Par-EPCA that may cause attention weight redundancy.
Figure 5: The performance comparsion of four SCFM implementations on the Fundus-PM dataset and the Isee-PA dataset.
@@ 5.1.3 The impact of three channel dependency modeling methods
n the top of our pyramid pooling (PP), we compare linear-SCFM with a multi-layer perception (MLP) with two fully connected layers (used in SPA and SE), Shared-MLP (employed in CBAM), and channel-interaction fully connected (CIC) layer (employed in ECA). To construct MLP based on the PP, we concatenate the multi-scale context features along the channel axis and follow the default settings of SPA and SE. Table 2 presents the PM classification results of four channel dependency modeling methods. We see that linear-SCFM outperforms MLP, share-MLP, and CIC despite its simplicity, highlighting the advantage of exploiting multi-scale context features in a channel-independent method.
@@ 5.3 Comparisons with state-of-the-art deep neural networks
Table 4 lists the PM recognition results of EPCA-Net and SOTA deep neural networks (e.g., RIR-Net, DR-Net, and MSCA-Net) on the Fundus-PM dataset and the WM-PM dataset. Note that our EPCA-Net gets the best PM recognition performance among all methods on five evaluation measures. For example, on the PM-Fundus dataset, EPCA-Net obtains over 1% and 2% gains of accuracy and kappa through comparisons to RIR-Net, MSA-Net, and the method proposed by Fu et al. fu2023automatic . Noticeably, compared with ViT, EPCA-Net obtains absolute over 2.85% gains of accuracy, 4.27 gains of sensitivity, 3.33% gains of F1, and 6.85% gains of kappa, respectively. As for the WM-PM dataset, EPCA-Net achieves above absolute 3.58% gains of five evaluation measures through comparisons to DR-Net, MSCA-Net, the method proposed by Fu et al. fu2023automatic , and FIT-Net. Remarkably, our EPCA-Net significantly outperforms SPA-ResNet50 and SE-ResNet50 by 4.77% of accuracy, 4.76% of sensitivity, 5.8% of F1, 11.59% of kappa, and 5.82% of AUC respectively. The above results verify that our method can effectively utilize the global-local pathology distribution prior information to enhance PM recognition performance, keeping consistent with our expectations.
@@
ACC Sen F1 Kappa AUC ACC Sen F1 Kappa AUC
RIR-Net ZHANG2022102499 96.47 96.03 96.03 92.07 98.43 88.10 84.13 84.13 68.25 79.90
DR-Net guo2020dense 96.88 96.13 96.47 92.93 99.12 85.71 80.95 80.95 61.90 80.54
MSCA-Net fu2023rmca 96.88 96.44 96.49 92.98 99.30 84.52 80.16 79.68 59.38 82.41
MSA-Net 9363019 90.64 88.00 89.14 78.31 95.58 83.33 79.37 78.45 56.92 75.39
Fu et al. fu2023automatic 96.20 95.53 95.71 91.42 99.10 84.52 80.16 79.68 59.38 79.40
FIT-Net chen2023fit 96.88 96.14 96.47 92.93 98.90 84.52 76.98 78.30 56.67 76.35
@@
ViT dosovitskiy2020image 94.71 93.09 93.93 87.87 98.12 84.52 81.75 80.27 60.61 80.84
SWiT liu2021swin 95.25 84.81 94.68 89.35 99.02 75.00 50.00 42.86 10.00 70.56
Res-MLP touvron2021resmlp 95.52 93.80 94.80 89.69 99.42 79.76 59.52 60.06 26.09 68.14
PVT wang2021pvtv2 96.34 95.12 95.82 91.64 99.51 83.33 68.25 71.59 45.10 81.37
@@
Table 5 presents the classification results of our method and SOTA deep neural networks on the Isee-PA dataset. We see that compared with CBAM-ResNet50, SPANet, RCRNet, ViT, and Res-MLP, EPCA-Net achieves 3.48% gains of ACC, F1, and kappa values. Remarkably, EPCA-Net outperforms MSCA-Net and DR-Net by over 3% and 13% in four evaluation metrics accordingly. Table 6 lists the glaucoma classification results of our EPCA-Net and state-of-the-art methods on the LAG dataset. It can be observed that EPCA-Net also achieves the best classification results on five evaluation measures. For example, EPCA-Net outperforms ViT and Swin-T by above absolute 3.6% in accuracy and 4.2% in sensitivity. Compared with RIR-Net and FIT-Net, our proposed method obtains absolute above 3.3% gains of accuracy. The results of Table 5 and Table 6 demonstrate the generalization ability of our method by dynamically adjusting the relative importance of multi-scale context features.
@@
Method ACC Sen F1 Kappa
RIR-Net ZHANG2022102499 75.94 74.24 75.33 62.59
DR-Net guo2020dense 63.26 62.46 62.63 43.72
MSCA-Net fu2023rmca 73.99 73.26 73.59 60.08
MSA-Net 9363019 76.63 76.21 76.47 64.22
Fu et al. fu2023automatic 76.50 75.13 76.13 63.58
FIT-Net chen2023fit 76.91 75.99 76.55 64.46
@@
ViT dosovitskiy2020image 68.29 66.00 65.92 50.55
Swin-T liu2021swin 76.08 75.54 95.75 63.37
Res-MLP touvron2021resmlp 66.48 63.87 63.33 47.51
@@
Method ACC Sen F1 Kappa AUC
CNN li2019large 89.2 90.6 95.60
Li et al. li2016integrating 89.7 91.4 96.00
RIR-Net ZHANG2022102499 93.11 92.75 92.50 85.00 97.60
DR-Net guo2020dense 93.42 92.66 92.77 85.54 97.83
MSCA-Net fu2023rmca 93.00 92.01 92.29 84.58 97.92
MSA-Net 9363019 88.99 87.92 87.94 75.88 94.77
Fu et al. fu2023automatic 92.90 91.86 92.17 84.34 97.97
FIT-Net chen2023fit 92.28 91.39 91.52 83.04 97.59
@@
ViT dosovitskiy2020image 89.92 89.56 89.09 78.18 96.42
Swin-T  liu2021swin 91.80 90.95 91.17 82.34 96.91
ResMLPtouvron2021resmlp 91.98 90.62 91.11 82.22 97.46
@@ 5.5.2 Weight visualization of multi-scale context features
Refer to caption (a) The relative weight statistics distributions of multi-scale context features at Res-EPCA_2_2
Refer to caption (c) The relative weight statistics distributions of multi-scale context features at Res-EPCA__\__4_6
Refer to caption (c) The relative weight statistics distributions of multi-scale context features at Res-EPCA_5_3

### v3

@@ 4.1 Datasets
PM-Fundus: To the best of our knowledge, the most existing datasets for PM recognition are private, lacking public PM recognition datasets. In order to prompt the development of PM recognition task, we collect fundus images of PM from publicly available PLAM dataset, ODIR dataset and RFMiD dataset [17] to construct a new dataset named PM-fundus, which will be released as a benchmark for automatic PM recognition task. The PM-fundus dataset contains 818 images of PM and 1636 images of normal. We split the PM-Fundus dataset into two disjoint sets: training and testing. The training set contains 572 images of PM and 1145 images of normal, and the testing set contains 246 images of PM and 491 images of normal, which will be released in the future.
@@
LAG: It is a publicly available fundus image dataset of glaucoma [50], which was collected by Beijing Tongren Hospital. The dataset has 4,854 images: 1,711 images of glaucoma and 3,143 images of non-glaucoma. We also split it into three disjoint subsets: training (2,911), validation (971), and testing (972).
@@ 4.2 Baselines
In this paper, we use the following attention methods to verify the effectiveness of our EPCA module: SE, SRM, ECA, SPA, external attention (EA) [51], coordination attention (CA) [52], and convolutional bottleneck attention module (CBAM) [53]. In addition, recent deep neural networks, such as vision transformer (ViT) [54], Res-MLP  [55], swin transformer (Swin-T)  [56] and pyramid vision transformer (PVT)  [57], and deep neural networks specifically designed for medical image analysis like DR-Net [58], MSCA-Net [59], MSA-Net [35], FIT-Net [32], RIR-Net [60], which are also adopted for comparison.
@@ 4.3 Implementation details
In this paper, we implement our EPCA-Net and other comparable deep neural networks with the Pytorch tool by using pre-trained backbone networks (such as ResNet18 and ResNet50) with traditional finetuning paradigm. We run all experiments on a server with an NVIDIA TITAN V GPU (11GB RAM). The SGD optimizer is utilized to optimize the deep networks with default settings (a weight decay of 1e-5 and a momentum of 0.9). We set the batch size and training epochs to 32 and 100, respectively. The initial learning rate (lr) is set to 0.0015, and the cosine annealing warm restart method [61] is applied to adjust lr. The classical data augmentation methods like flipping and rotation ($-10^{\circ}$-$10^{\circ}$) are used for data augmentation in the training.
@@ 4.4 Evaluation Metrics
Following existing work [33], we adopt the five evaluation metrics to verify the general performance of methods: accuracy (ACC), sensitivity (Sen), F1, kappa coefficient value, and the area under the ROC curve (AUC). In particular, AUC is applied to evaluate the binary classification performance of methods.
@@ 5.1.1 The impact of receptive field size number selection in pyramid pooling
As shown in Eq. 1, the receptive field size number selection in the pyramid pooling operator may significantly impact the EPCA module. Therefore, we train them with EPCA using Linear-SCFM implementation on the Fundus-PM and Isee-PA datasets. Here, we adopt the pre-trained ResNet18 and ResNet50 as the backbones. As shown in Fig. 2(b), this paper places our EPCA module behind the residual block for constructing the Res-EPCA module, where the parameters of the residual block loaded from pre-trained ResNets, and we only randomly initialize parameters of the EPCA module.
Table 2 lists the classification results of our EPCA with six different receptive field size selection settings. We see that multi-scale context features obtained by spatial average pooling operations with different receptive field sizes affect the performance of EPCA, demonstrating that multi-scale context features have varying significance levels. When pyramid pooling with multiple receptive filed size setting $\{1,3\}$, our EPCA achieves the best accuracies on two datasets through two backbones, which is used for the following experiments. Moreover, the EPCA module based on ResNet50 performs better than it based on ResNet18; thus, ResNet50 is adopted as the backbone for the following ablation experiments.
Table 2: Performance comparisons of receptive field size number selections in pyramid pooling based on the EPCA module
Backbone Receptive field size setting ACC (%)
PM-Fundus Isee-PA
ResNet18 $\left\{1\right\}$ 96.07 74.27
$\left\{1,2\right\}$ 96.61 74.83
$\left\{1,3\right\}$ 96.88 76.50
$\left\{1,4\right\}$ 96.07 76.35
$\left\{1,2,3\right\}$ 95.93 74.69
$\left\{1,2,4\right\}$ 95.79 76.36
ResNet50 $\left\{1\right\}$ 96.47 74.41
$\left\{1,2\right\}$ 96.20 76.63
$\left\{1,3\right\}$ 97.56 77.19
$\left\{1,4\right\}$ 96.93 75.80
$\left\{1,2,3\right\}$ 96.47 75.94
$\left\{1,2,4\right\}$ 95.93 76.36
@@ 5.1.2 The impact of different SCFM implementations
Fig. 6 presents the classification results of the EPCA module with four different SCFM implementations on the Fundus-PM (green color) dataset and the Isee-PA (blue color) dataset. It can be seen that EPCA module with Linear-SCFM and Hierarchical-SCFM achieves similar performance and outperforms it with Dropout-SCFM and Parallel-SCFM. The following reasons can explain the results: 1) The Dropout technique randomly selects context features, which makes it challenging to evaluate whether the selected context features are helpful while the rest of the context features are useless. 2) Although multi-scale context features with varying significance levels are extracted from different receptive sizes based on spatial pyramid average pooling, every receptive filed size produces attention weights in the Par-EPCA that may cause attention weight redundancy.
Figure 6: The performance comparsion of four SCFM implementations on the Fundus-PM dataset and the Isee-PA dataset by using accuracy (ACC) as the evaluation measure.
@@ 5.1.3 The impact of three channel dependency modeling methods
On the top of our pyramid pooling (PP), we compare linear-SCFM with a multi-layer perception (MLP) include two fully connected layers (used in SPA and SE), Shared-MLP (employed in CBAM), and channel-interaction fully connected (CIC) layer (employed in ECA). To construct MLP based on the PP, we concatenate the multi-scale context features along the channel axis and follow the default settings of SPA and SE. Table 3 presents the PM classification results of four channel dependency modeling methods. Despite its simplicity, we see that linear-SCFM outperforms MLP, share-MLP, and CIC, highlighting the advantage of exploiting multi-scale context features in a channel-independent manner.
@@ 5.3 Comparisons with state-of-the-art deep neural networks
Table 5 lists the PM recognition results of EPCA-Net and SOTA deep neural networks (e.g., RIR-Net, DR-Net, and MSCA-Net) on the Fundus-PM dataset and the WM-PM dataset. Note that our EPCA-Net gets the best PM recognition performance among all methods on five evaluation measures. For example, on the PM-Fundus dataset, EPCA-Net obtains over 1% and 2% gains of accuracy and kappa through comparisons to RIR-Net, MSA-Net, and the method proposed by Fu et al. [38]. Noticeably, compared with ViT, EPCA-Net obtains absolute over 2.85% gains of accuracy, 4.27% gains of sensitivity, 3.33% gains of F1, and 6.65% gains of kappa, respectively. As for the WM-PM dataset, EPCA-Net achieves above absolute 3.58% gains of five evaluation measures through comparisons to DR-Net, MSCA-Net, the method proposed by Fu et al. [38], and FIT-Net. Remarkably, our EPCA-Net significantly outperforms SPA-ResNet50 and SE-ResNet50 by 4.77% of accuracy, 4.76% of sensitivity, 5.8% of F1, 11.59% of kappa, and 5.82% of AUC respectively. The above results verify that our method can effectively utilize the global-local pathology distribution prior information to enhance PM recognition performance, keeping consistent with our expectations.
@@
ACC (%) Sen (%) F1 (%) Kappa (%) AUC (%) ACC (%) Sen (%) F1 (%) Kappa (%) AUC (%)
RIR-Net [60] 96.47 96.03 96.03 92.07 98.43 88.10 84.13 84.13 68.25 79.90
DR-Net [58] 96.88 96.13 96.47 92.93 99.12 85.71 80.95 80.95 61.90 80.54
MSCA-Net [59] 96.88 96.44 96.49 92.98 99.30 84.52 80.16 79.68 59.38 82.41
MSA-Net [35] 90.64 88.00 89.14 78.31 95.58 83.33 79.37 78.45 56.92 75.39
Fu et al. [38] 96.20 95.53 95.71 91.42 99.10 84.52 80.16 79.68 59.38 79.40
FIT-Net [32] 96.88 96.14 96.47 92.93 98.90 84.52 76.98 78.30 56.67 76.35
@@
ViT [54] 94.71 93.09 93.93 87.87 98.12 84.52 81.75 80.27 60.61 80.84
SWiT [56] 95.25 84.81 94.68 89.35 99.02 75.00 50.00 42.86 10.00 70.56
Res-MLP [55] 95.52 93.80 94.80 89.69 99.42 79.76 59.52 60.06 26.09 68.14
PVT [57] 96.34 95.12 95.82 91.64 99.51 83.33 68.25 71.59 45.10 81.37
@@
Table 6 presents the classification results of our method and SOTA deep neural networks on the Isee-PA dataset. We see that compared with CBAM-ResNet50, SPANet, RCRNet, ViT, and Res-MLP, EPCA-Net achieves 3.48% gains of ACC, F1, and kappa values. Remarkably, EPCA-Net outperforms MSCA-Net and DR-Net by over 3.16% and 13.93% in four evaluation metrics accordingly. Table 7 lists the glaucoma classification results of our EPCA-Net and state-of-the-art methods on the LAG dataset. It can be observed that EPCA-Net also achieves the best classification results on five evaluation measures. For example, EPCA-Net outperforms ViT and Swin-T by above absolute 3.78% of accuracy and 4.24% of sensitivity, respectively. Compared with RIR-Net and FIT-Net, our proposed method obtains absolute above 2.27% gains of accuracy. The results of Table 6 and Table 7 demonstrate the generalization ability of our method by dynamically adjusting the relative importance of multi-scale context features.
@@
Method ACC (%) Sen (%) F1 (%) Kappa (%)
RIR-Net [60] 75.94 74.24 75.33 62.59
DR-Net [58] 63.26 62.46 62.63 43.72
MSCA-Net [59] 73.99 73.26 73.59 60.08
MSA-Net [35] 76.63 76.21 76.47 64.22
Fu et al. [38] 76.50 75.13 76.13 63.58
FIT-Net [32] 76.91 75.99 76.55 64.46
@@
ViT [54] 68.29 66.00 65.92 50.55
Swin-T [56] 76.08 75.54 95.75 63.37
Res-MLP [55] 66.48 63.87 63.33 47.51
@@
Method ACC (%) Sen (%) F1 (%) Kappa (%) AUC (%)
CNN [50] 89.2 90.6 95.60
Li et al. [62] 89.7 91.4 96.00
RIR-Net [60] 93.11 92.75 92.50 85.00 97.60
DR-Net [58] 93.42 92.66 92.77 85.54 97.83
MSCA-Net [59] 93.00 92.01 92.29 84.58 97.92
MSA-Net [35] 88.99 87.92 87.94 75.88 94.77
Fu et al. [38] 92.90 91.86 92.17 84.34 97.97
FIT-Net [32] 92.28 91.39 91.52 83.04 97.59
@@
ViT [54] 89.92 89.56 89.09 78.18 96.42
Swin-T  [56] 91.80 90.95 91.17 82.34 96.91
ResMLP[55] 91.98 90.62 91.11 82.22 97.46
@@
Table 8 lists the five-fold cross-validation results of our EPCA-Net and four other representative deep neural networks on the PM-Fundus dataset. Here, we split the PM-Fundus dataset into five disjoint sets; at each time, four sets are used for training (one set is adopted as the validation set for selecting the best model), and a set is used for testing. It can be seen that our method performs better than comparable deep neural networks and is also more robust than them.
Table 9 presents the statistical significance analysis results among EPCA-Net and representative deep neural networks with Student’s t-test (t-test) method. We find significant differences between our proposed EPCA-Net and comparable deep neural networks (p-value$<$0.05), suggesting that learned feature representations are different.
Table 8: Five-fold cross-validation among EPCA-Net and representative deep neural networks method on PM-fundus dataset
Method ACC (%) Sen (%) F1 (%) Kappa (%) AUC (%)
RIR-Net 79.01 $\pm$ 8.43 72.61 $\pm$ 23.36 68.23 $\pm$ 28.92 61.43 $\pm$ 29.91 70.74 $\pm$ 7.21
FIT-Net 88.95 $\pm$ 5.41 82.41 $\pm$ 17.14 84.42 $\pm$ 15.92 70.21 $\pm$ 24.56 77.40 $\pm$ 4.36
SE-Net 88.12 $\pm$ 11.11 81.42 $\pm$ 15.61 80.21 $\pm$ 22.65 69.22 $\pm$31.25 78.48 $\pm$ 13.24
SPA-Net 85.81 $\pm$ 8.67 82.83 $\pm$ 13.39 83.11 $\pm$ 12.00 67.12 $\pm$ 24.61 86.28 $\pm$ 8.13
EPCA-Net 90.42 $\pm$ 4.62 83.61 $\pm$ 13.07 84.83 $\pm$ 11.90 70.81 $\pm$ 23.64 87.80 $\pm$ 3.97
Table 9: Significant analysis among EPCA-Net and representative deep neural networks with t-test method on PM-fundus dataset
Comparison p-value
EPCA-Net $Vs.$ RIR-Net 0.004
EPCA-Net $Vs.$ FIT-Net 0.01
EPCA-Net $Vs.$ SE-Net 0.0059
EPCA-Net $Vs.$ SPA-Net 0.0082
@@ 5.5.2 Weight visualization of multi-scale context features
(a) The relative weight statistics distributions of multi-scale context features at Res-EPCA_2_2
(c) The relative weight statistics distributions of multi-scale context features at Res-EPCA__\__4_6
(c) The relative weight statistics distributions of multi-scale context features at Res-EPCA_5_3